OpenStudy (anonymous):

i need help with factor by grouping. ac - bc + ad - bd xy - 3x - 8y + 24 4x^2 + 20x + 5y +xy 48x^2y + 20xz +12xy + 5z

hero (hero):

So for the first one: ac - bc + ad - bd = c(a-b)+d(a-b) = (a-b)(c+d) Notice that in order to factor by grouping, there has to be a term common to both both expressions